# CONSENT: A Negotiation Framework for Leveraging User Flexibility in Vehicle-to-Building Charging under Uncertainty

### v1

@@ 7. Results
Baseline Policy Performance. The real-world baseline of uncoordinated free charging (Policy 2) (al2021slow) raises the building’s monthly cost by over 6%. A smart scheduling algorithm (Policy 3) (sen2025; liu2025) mitigates this increase but still costs more than the no EV reference. Introducing a simple, fixed price for users (Policy 4) based on the utility’s peak energy pricing (svp) further reduces the building’s cost, bringing it nearly to the ideal no EV baseline. Finally, using our utility-based, demand-aware user pricing (Policy 5) enables the building to achieve its lowest operational cost among all policies. However, this purely operator-centric optimization leads to the highest per-unit user cost, leading to negotiation-based approaches.
Negotiation Framework Performance. Building on these baselines, our first negotiation-based comparison is Policy 5, which implements a state-of-the-art, menu-based approach consistent with prior work (ghosh2017menu; bae2021inducing). While this method reduces user costs, it exhibits a high rejection rate of nearly 35%, indicating misaligned incentives.
Our proposed framework, CONSENT (Policy 6), strikes the best overall balance between operator and user objectives, and offers a 12% lower user cost at \$0.138/kWh and substantially reduces the rejection rate from 34.89% to 21.99%, compared to Menu-based Negotiation. This user cost is over 61% lower than average public charging rates in the United States (ev_price). Simultaneously, the framework reduces building operator costs by 5.83% compared to uncoordinated charging and by nearly 3.5% compared to an optimized smart-charging policy with free user access (Policy 3).
@@
Figure 2. Comparison of computational runtime per decision across varying EV arrival intensities.
@@
Menu-based Negotiation (bae2021inducing; ghosh2017menu)
@@
Table 2. Sensitivity of negotiation outcomes to different user flexibility profiles where we compare our framework, CONSENT, against a menu-based and an ablated baseline.
@@
Table 3. Performance across utility-sharing ratios $\alpha$ from user cost equation (3) (100% to 10%): lower utility sharing lowers building cost but raises user costs and simulated rejections, highlighting trade-offs in operational savings and user acceptance.
@@ Appendix A Appendix
Table 4. Charger Assignment and tiebreaker comparisons with an oracle, optimal policy (MILP). Assigning to Bidirectional chargers first and then breaking ties by assigning them to the EV that departs later results in the lowest bill. Lower is better.
@@
Table 5. Key Input Variables
@@ Appendix B Charging Policy
Table 6. Key Decision Variables
@@
(6) $\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v}\setminus\{t_{\text{arr}}^{v}\}:\quad a^{v,c}_{t-1}-a^{v,c}_{t}=0$
@@
(7) $\footnotesize\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v},\ \forall\delta\in\Delta:\quad\begin{cases}e^{v}_{t}\geq e^{v,\delta}_{\min}\cdot y^{v,\delta}_{t}-(1-y^{v,\delta}_{t})\\ e^{v}_{t}\leq e^{v,\delta}_{\max}\cdot y^{v,\delta}_{t}+M_{1}(1-y^{v,\delta}_{t})\end{cases}$
@@
(8) $\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}:\quad\sum_{\delta\in\Delta}y^{v,\delta}_{t}=1$
@@
(9) $x^{v,c,\delta}_{t}\leq a^{v,c}_{t};\quad x^{v,c,\delta}_{t}\leq y^{v,\delta}_{t};\quad x^{v,c,\delta}_{t}\geq a^{v,c}_{t}+y^{v,\delta}_{t}-1$
(10) $\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:r^{v}_{t}\leq\eta\sum_{c\in\mathcal{C}}\sum_{\delta\in\Delta}r^{v,\delta}_{\max}\cdot x^{v,c,\delta}_{t}$
@@
(11) $\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:\quad r^{v}_{t}\geq\eta\sum_{c\in\mathcal{C}}c_{\min}\cdot a_{v,c,t}$
@@
(12) $\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:\quad e^{v}_{t_{n}}=\begin{cases}e^{v}_{t_{arr}}+r^{v}_{t_{n}},&\text{if }n=0\\[8.0pt] e^{v}_{t_{n-1}}+r^{v}_{t_{n}},&\text{otherwise}\end{cases}$
@@
(14) $\forall t\in\mathcal{T}:\quad g_{t}=\sum_{v\in\mathcal{V}}w_{t}\left(c^{v}_{t}+{p^{b}_{t}}/{\tau}\right)$
@@
(16) $\forall t\in\mathcal{T}:\quad p_{t}=p^{b}_{t}+\frac{1}{\tau}\sum_{v\in\mathcal{V}}r^{v}_{t}$
(17) $\forall t\in\mathcal{T}:\quad p^{\max}\geq p_{t}$
@@
(18) $\begin{split}\textstyle\min\sum_{t\in\mathcal{T}}g_{t}+K_{\mathrm{DC}}\cdot p^{\max}+K_{\mathrm{SOC}}\cdot\sum_{v\in\mathcal{V}}z_{v}\\ +K_{batt}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q^{v}_{t}\end{split}$
@@
B.1. Online Control
In real-world deployment, we consider the problem of optimizing EV charging decisions in a dynamic environment, where the static optimization method described earlier becomes inapplicable due to the absence of oracle inputs. To handle this online decision-making process under uncertainty, we propose a Monte Carlo Model Predictive Control (MC-MPC) framework, which belongs to the family of stochastic Model Predictive Control (SMPC) methods [MA2016]. This framework extends our static optimization approach to an online setting by leveraging predictive models to simulate future scenarios.
@@ B.1. Online Control
(19) $\min_{\begin{subarray}{c}A_{t},\,\{A_{t+1}^{f},\dots,A_{H(t)}^{f}\}_{f\in\mathcal{F}}\end{subarray}}\;\mathbb{E}_{f\in\mathcal{F}}\Bigg[$ $\sum_{k=t}^{H(t)}g_{k,f}+w^{d}p^{\max}_{f}$
@@
(20) $\tilde{r}_{t}^{v}=\begin{cases}r_{t}^{v}+\min\left(\frac{P_{\text{diff}}}{|\mathcal{V}_{t}|}\tau,\ e_{\text{req}}^{v},\ \phi(v)_{{max}}\right),&\text{if }\phi(v)_{{min}}=0\\ r_{t}^{v}+\min\left(\frac{P_{\text{diff}}}{|\mathcal{V}_{t}|}\tau,\ e_{\max}^{v},\ \phi(v)_{\max}\right),&\text{if }\phi(v)_{{min}}<0\end{cases}$
@@ Appendix C Optimization Framework for Negotiation
    (21) $\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v},t>t^{v}_{\text{arr}}:$
    (22) $\quad a^{v,cp}_{t-1}\leq a^{v,cp}_{t}+\lambda^{v}_{t}$
  3. (3)
@@
    (23) $\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v}:\quad a_{v,c,t}\leq 1-\sum_{t=t^{v}_{\text{arr}}}^{t}\lambda^{v}_{t}$
  4. (4)
@@
    (24) $\forall v\in\mathcal{V}:\quad t_{\text{dep}}^{v}=\sum_{t\in\mathcal{T}^{v}}t\cdot\lambda_{v,t}$
  7. (7)
@@
    (25) $\forall v\in\mathcal{V}:\quad\sum_{s\in\mathcal{T}^{v}}\lambda^{v}_{t}=1$
  8. (8)
@@
    (26) $\forall v\in\mathcal{V}:\quad\sigma_{soc}^{v}=e_{\text{req}}^{v}-e_{\text{dep}}^{v}$
    (27) $\forall v\in\mathcal{V}:\quad\sigma_{dep}^{v}=t_{\text{dep}}^{v}-t_{\text{req}}^{v}$
  9. (9)
@@
(28) $\begin{split}\textstyle\min\sum_{t\in\mathcal{T}}g_{t}+K_{\mathrm{DC}}\cdot p_{\max}+K_{\mathrm{SOC}}\cdot\sum_{v\in\mathcal{V}}z_{v}\\ +K_{batt}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q^{v}_{t}\end{split}$
@@ Appendix D Proof of Theorems
(29) $U(\bar{\theta}_{v})=J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bar{\theta}_{v},\bm{\hat{\theta}}_{\setminus v}).$
@@
(30) $U(\theta_{v})-U(\bar{\theta}_{v})=J^{\pi}(\theta_{\setminus v},\bar{\theta}_{v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi}(\theta_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v}).$
@@
(33) $J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})-\zeta^{v}\leq J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}).$
@@
(34) $\zeta^{v}\leq J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})=U(\theta_{v}).$
@@
(36) $J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})+U(\theta_{v})=J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}).$
@@ Appendix E User Segmentation from Survey
E.1. Flexibility limits for $L$ Negotiation Options
@@
2233445566770.480.480.490.490.50.50.510.510.520.520.530.530.540.540.550.550.560.560.570.570.580.580.590.590.60.6Optimal k=6k=6Most SuitableNumber of ClustersSilhouette ScoreNegotiation Options
Figure 3. Silhouette method shows $k=6$ as the optimal choice (highest score). However, $k=3$ is numerically very close, and is most suitable as users might get overwhelmed with more options.
223344556677224466881010121214141616181820202222242426262828Elbow PointNumber of ClustersWCSSNegotiation Options
Figure 4. Elbow method shows a sharp decrease until $k=3$, indicating the elbow point.
Figure 5. Clustered user deviation limits data provides Negotiation Options and flexibility limits
@@
Figure 6. Cluster selection and segmentation of user types. Left: The Elbow method plot shows inertia (WCSS) against the number of clusters, with an identified elbow at $k=4$. Center: Silhouette scores for different cluster counts, indicating the best silhouette value at $k=4$. Right: Final clustering of survey respondents into four user types based on SoC sensitivity ($w_{1}$) and time sensitivity ($w_{2}$), with cluster centroids marked by black crosses. Each color represents a distinct user type.
E.2. Cluster Characterization and Representative Points
@@
Table 7. Negotiation Options (Flexibility Analysis) Summary
@@
Table 8. User Type Characteristics: Survey input points
@@
E.3. User Choice Modeling

### v3

@@ 7 Results
Baseline Policy Performance. The real-world baseline of uncoordinated free charging (Policy 2) [3] raises the building’s monthly cost by over 6%. A smart scheduling algorithm (Policy 3) [19, 13] mitigates this increase but still costs more than the no EV reference. Introducing a simple, fixed price for users (Policy 4) based on the utility’s peak energy pricing [22] further reduces the building’s cost, bringing it nearly to the ideal no EV baseline. Finally, using our utility-based, demand-aware user pricing (Policy 5) enables the building to achieve its lowest operational cost among all policies. However, this purely operator-centric optimization leads to the highest per-unit user cost, leading to negotiation-based approaches.
Negotiation Framework Performance. Building on these baselines, our first negotiation-based comparison is Policy 5, which implements a state-of-the-art, menu-based approach consistent with prior work [9, 5]. While this method reduces user costs, it exhibits a high rejection rate of nearly 35%, indicating misaligned incentives.
Our proposed framework, CONSENT (Policy 6), strikes the best overall balance between operator and user objectives, and offers a 12% lower user cost at \$0.138/kWh and substantially reduces the rejection rate from 34.89% to 21.99%, compared to Menu-based Negotiation. This user cost is over 61% lower than average public charging rates in the United States [1]. Simultaneously, the framework reduces building operator costs by 5.83% compared to uncoordinated charging and by nearly 3.5% compared to an optimized smart-charging policy with free user access (Policy 3).
@@
Figure 3: Comparison of computational runtime per decision across varying EV arrival intensities.
@@
Menu-based Negotiation [5, 9]
@@
Table 2: Sensitivity of negotiation outcomes to different user flexibility profiles where we compare our framework, CONSENT, against a menu-based and an ablated baseline.
@@
Table 3: Performance across utility-sharing ratios $\alpha$ from user cost equation (3) (100% to 10%): lower utility sharing lowers building cost but raises user costs and simulated rejections, highlighting trade-offs in operational savings and user acceptance.
@@ 8 Conclusion
9 Acknowledgements
This material is based upon work sponsored by Nissan Advanced Technology Center-Silicon Valley. Results presented in this paper were obtained using the Chameleon Testbed supported by the National Science Foundation. Any opinions, findings, conclusions, or recommendations expressed in this material are those of the authors and do not necessarily reflect the views of Nissan.
@@ Appendix A Appendix
Table 4: Charger Assignment and tiebreaker comparisons with an oracle, optimal policy (MILP). Assigning to Bidirectional chargers first and then breaking ties by assigning them to the EV that departs later results in the lowest bill. Lower is better.
@@
Table 5: Key Input Variables
@@ Appendix B Charging Policy
Table 6: Key Decision Variables
@@
$\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v}\setminus\{t_{\text{arr}}^{v}\}:\quad a^{v,c}_{t-1}-a^{v,c}_{t}=0$ (6)
@@
$\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v},\ \forall\delta\in\Delta:\quad\begin{cases}e^{v}_{t}\geq e^{v,\delta}_{\min}\cdot y^{v,\delta}_{t}-(1-y^{v,\delta}_{t})\\ e^{v}_{t}\leq e^{v,\delta}_{\max}\cdot y^{v,\delta}_{t}+M_{1}(1-y^{v,\delta}_{t})\end{cases}$ (7)
@@
$\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}:\quad\sum_{\delta\in\Delta}y^{v,\delta}_{t}=1$ (8)
@@
$x^{v,c,\delta}_{t}\leq a^{v,c}_{t};\quad x^{v,c,\delta}_{t}\leq y^{v,\delta}_{t};\quad x^{v,c,\delta}_{t}\geq a^{v,c}_{t}+y^{v,\delta}_{t}-1$ (9)
$\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:r^{v}_{t}\leq\eta\sum_{c\in\mathcal{C}}\sum_{\delta\in\Delta}r^{v,\delta}_{\max}\cdot x^{v,c,\delta}_{t}$ (10)
@@
$\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:\quad r^{v}_{t}\geq\eta\sum_{c\in\mathcal{C}}c_{\min}\cdot a_{v,c,t}$ (11)
@@
$\forall v\in\mathcal{V},\ \forall t\in\mathcal{T}^{v}:\quad e^{v}_{t_{n}}=\begin{cases}e^{v}_{t_{arr}}+r^{v}_{t_{n}},&\text{if }n=0\\[8.0pt] e^{v}_{t_{n-1}}+r^{v}_{t_{n}},&\text{otherwise}\end{cases}$ (12)
@@
$\forall t\in\mathcal{T}:\quad g_{t}=\sum_{v\in\mathcal{V}}w_{t}\left(c^{v}_{t}+{p^{b}_{t}}/{\tau}\right)$ (14)
@@
$\forall t\in\mathcal{T}:\quad p_{t}=p^{b}_{t}+\frac{1}{\tau}\sum_{v\in\mathcal{V}}r^{v}_{t}$ (16)
$\forall t\in\mathcal{T}:\quad p^{\max}\geq p_{t}$ (17)
@@
$\begin{split}\textstyle\min\sum_{t\in\mathcal{T}}g_{t}+K_{\mathrm{DC}}\cdot p^{\max}+K_{\mathrm{SOC}}\cdot\sum_{v\in\mathcal{V}}z_{v}\\ +K_{batt}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q^{v}_{t}\end{split}$ (18)
@@
B.1 Online Control
In real-world deployment, we consider the problem of optimizing EV charging decisions in a dynamic environment, where the static optimization method described earlier becomes inapplicable due to the absence of oracle inputs. To handle this online decision-making process under uncertainty, we propose a Monte Carlo Model Predictive Control (MC-MPC) framework, which belongs to the family of stochastic Model Predictive Control (SMPC) methods [16]. This framework extends our static optimization approach to an online setting by leveraging predictive models to simulate future scenarios.
@@ B.1 Online Control
$\min_{\begin{subarray}{c}A_{t},\,\{A_{t+1}^{f},\dots,A_{H(t)}^{f}\}_{f\in\mathcal{F}}\end{subarray}}\;\mathbb{E}_{f\in\mathcal{F}}\Bigg[$ $\sum_{k=t}^{H(t)}g_{k,f}+w^{d}p^{\max}_{f}$ (19)
@@
$\tilde{r}_{t}^{v}=\begin{cases}r_{t}^{v}+\min\left(\frac{P_{\text{diff}}}{|\mathcal{V}_{t}|}\tau,\ e_{\text{req}}^{v},\ \phi(v)_{{max}}\right),&\text{if }\phi(v)_{{min}}=0\\ r_{t}^{v}+\min\left(\frac{P_{\text{diff}}}{|\mathcal{V}_{t}|}\tau,\ e_{\max}^{v},\ \phi(v)_{\max}\right),&\text{if }\phi(v)_{{min}}<0\end{cases}$ (20)
@@ Appendix C Optimization Framework for Negotiation
    $\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v},t>t^{v}_{\text{arr}}:$ (21)
    $\quad a^{v,cp}_{t-1}\leq a^{v,cp}_{t}+\lambda^{v}_{t}$ (22)
  3. 3.
@@
    $\forall v\in\mathcal{V},\ \forall c\in\mathcal{C},\ \forall t\in\mathcal{T}^{v}:\quad a_{v,c,t}\leq 1-\sum_{t=t^{v}_{\text{arr}}}^{t}\lambda^{v}_{t}$ (23)
  4. 4.
@@
    $\forall v\in\mathcal{V}:\quad t_{\text{dep}}^{v}=\sum_{t\in\mathcal{T}^{v}}t\cdot\lambda_{v,t}$ (24)
  7. 7.
@@
    $\forall v\in\mathcal{V}:\quad\sum_{s\in\mathcal{T}^{v}}\lambda^{v}_{t}=1$ (25)
  8. 8.
@@
    $\forall v\in\mathcal{V}:\quad\sigma_{soc}^{v}=e_{\text{req}}^{v}-e_{\text{dep}}^{v}$ (26)
    $\forall v\in\mathcal{V}:\quad\sigma_{dep}^{v}=t_{\text{dep}}^{v}-t_{\text{req}}^{v}$ (27)
  9. 9.
@@
$\begin{split}\textstyle\min\sum_{t\in\mathcal{T}}g_{t}+K_{\mathrm{DC}}\cdot p_{\max}+K_{\mathrm{SOC}}\cdot\sum_{v\in\mathcal{V}}z_{v}\\ +K_{batt}\sum_{v\in\mathcal{V}}\sum_{t\in\mathcal{T}^{v}}q^{v}_{t}\end{split}$ (28)
@@ Appendix D Proof of Theorems
$U(\bar{\theta}_{v})=J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bar{\theta}_{v},\bm{\hat{\theta}}_{\setminus v}).$ (29)
@@
$U(\theta_{v})-U(\bar{\theta}_{v})=J^{\pi}(\theta_{\setminus v},\bar{\theta}_{v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi}(\theta_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v}).$ (30)
@@
$J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})-\zeta^{v}\leq J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}).$ (33)
@@
$\zeta^{v}\leq J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v})-J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})=U(\theta_{v}).$ (34)
@@
$J^{\pi^{*}}(\bm{\theta}_{\setminus v},\theta_{v},\bm{\hat{\theta}}_{\setminus v})+U(\theta_{v})=J^{\pi^{*}}(\bm{\theta}_{\setminus v},\bm{\hat{\theta}}_{\setminus v}).$ (36)
@@ Appendix E User Segmentation from Survey
E.1 Flexibility limits for $L$ Negotiation Options
@@
2233445566770.480.480.50.50.520.520.540.540.560.560.580.580.60.6Optimal k=6k=6Most SuitableNumber of ClustersSilhouette ScoreNegotiation Options
Figure 4: Silhouette method shows $k=6$ as the optimal choice (highest score). However, $k=3$ is numerically very close, and is most suitable as users might get overwhelmed with more options.
223344556677551010151520202525Elbow PointNumber of ClustersWCSSNegotiation Options
Figure 5: Elbow method shows a sharp decrease until $k=3$, indicating the elbow point.
Figure 6: Clustered user deviation limits data provides Negotiation Options and flexibility limits
@@
Figure 7: Cluster selection and segmentation of user types. Left: The Elbow method plot shows inertia (WCSS) against the number of clusters, with an identified elbow at $k=4$. Center: Silhouette scores for different cluster counts, indicating the best silhouette value at $k=4$. Right: Final clustering of survey respondents into four user types based on SoC sensitivity ($w_{1}$) and time sensitivity ($w_{2}$), with cluster centroids marked by black crosses. Each color represents a distinct user type.
E.2 Cluster Characterization and Representative Points
@@
Table 7: Negotiation Options (Flexibility Analysis) Summary
@@
Table 8: User Type Characteristics: Survey input points
@@
E.3 User Choice Modeling